# Tackling Time-Series Forecasting Generalization via Mitigating Concept Drift

### v1

@@ 1 Introduction
Time-series forecasting finds applications in various real-world scenarios such as economics, urban computing, and epidemiology (Zhu & Shasha, 2002; Zheng et al., 2014; Deb et al., 2017; Mathis et al., 2024). These applications involve predicting future trends or events based on historical time-series data. For example, economists use forecasts to make financial and marketing plans, while sociologists use them to allocate resources and formulate policies for traffic or disease control.
@@ 2.2 Distribution Shift in Time Series
Temporal shift (also known as virtual shift (Tsymbal, 2004)) is the marginal probability distributions changing over time, while the conditional distributions are the same.
@@
Concept drift (also known as real concept drift (Gama et al., 2014)111(Gama et al., 2014) defines concept drift as both virtual shift and real concept drift. Our concept drift definition is consistent with the definition of real concept drift in (Gama et al., 2014).) is the conditional distributions changing over time, while the marginal probability distributions are the same.
@@ 4.1 Setup
Datasets. We conduct experiments using six time-series datasets as leveraged in (Liu et al., 2024a): The daily reported currency exchange rates (Exchange) (Lai et al., 2018); The weekly reported influenza-like illness patients (ILI) (Kamarthi et al., 2021); Two-hourly/minutely reported electricity transformer temperature (ETTh1/ETTh2 and ETTm1/ETTm2, respectively) (Zhou et al., 2021). We follow the established experimental setups and target variable selections in previous works(Wu et al., 2021; 2022; Nie et al., 2023; Liu et al., 2024b). Datasets such as Traffic (PeMS) (Zhao et al., 2017) and Weather (Wu et al., 2021) are excluded from our evaluations, as their time series exhibit near-stationary behavior, with only moderate distribution shift issues. Further details on the dataset differences are discussed in Appendix C.1.
@@
Forecasting Model Baselines: ShifTS is model-agnostic, we include six time-series forecasting models (referred to as ‘Model’ in Table 1 and 4), including: Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), Crossformer (Zhang & Yan, 2022), PatchTST (Nie et al., 2023), TimeMixer (Wang et al., 2024) and iTransformer (Liu et al., 2024b), which of the last two are the state-of-the-art (SOTA) forecasting model. These models are used to demonstrate that ShifTS consistently enhances forecasting accuracy across various models, including SOTA.
@@ 4.2 Performance Improvement across Base Forecasting Models
| Dataset | ILI | Exchange | ETTh1 | ETTh2 | ||||||||
| Method | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | ||||
| Base | ERM | 3.705 | 1.704 | 0.819 | 0.732 | 0.254 | 0.427 | 0.937 | 0.828 | |||
GroupDRO | 2.285 | 1.287 | 0.821 | 0.751 | 0.278 | 0.453 | 1.150 | 0.936 | |||
| IRM | 2.248 | 1.237 | 0.846 | 0.754 | 0.201 | 0.367 | 0.878 | 0.792 | ||||
| VREx | 2.285 | 1.286 | 0.821 | 0.742 | 0.314 | 0.486 | 1.142 | 0.938 | ||||
| EIIL | 2.036 | 1.159 | 0.822 | 0.749 | 0.212 | 0.433 | 1.122 | 0.930 | ||||
RevIN | 0.815 | 0.708 | 0.475 | 0.476 | 0.085 | 0.224 | 0.205 | 0.358 | |||
| N-S Trans. | 0.781 | 0.688 | 0.484 | 0.481 | 0.086 | 0.226 | 0.203 | 0.355 | ||||
| SAN | 0.757 | 0.715 | 0.415 | 0.453 | 0.088 | 0.225 | 0.199 | 0.348 | ||||
IRM+RevIN | 0.809 | 0.711 | 0.481 | 0.476 | 0.089 | 0.231 | 0.202 | 0.362 | |||
| EIIL+RevIN | 0.799 | 0.706 | 0.483 | 0.485 | 0.085 | 0.225 | 0.218 | 0.380 | ||||
| FOIL | 0.735 | 0.651 | 0.497 | 0.481 | 0.081 | 0.219 | 0.206 | 0.357 | ||||
| ShifTS (Ours) | 0.668 | 0.613 | 0.470 | 0.468 | 0.076 | 0.214 | 0.194 | 0.348 | ||||
@@ 4.3 Comparison with Distribution Shift Methods
The results highlight the advantages of ShifTS over existing distribution shift methods, achieving the highest average forecasting accuracy in 6 out of 8 evaluations, with the remaining 2 evaluations ranking second. Notably, as discussed in Section 3.2, we choose to use RevIN as it is one of the most popular yet simple and effective temporal shift methods. However, ShifTS is flexible and can integrate more advanced temporal shift methods to further enhance performance. While exploring these advanced temporal shift methods is beyond the scope of this work, we illustrate the potential benefits of such integration.
@@
For example, on the Exchange dataset, where SAN outperforms ShifTS, incorporating SAN in place of RevIN within ShifTS leads to even greater accuracy improvements. Detailed MSE values for these evaluations are provided in Table 3. Furthermore, the results underscore the importance of addressing concept drift using SAM when temporal shifts are effectively addressed.
@@ 6 Acknowledgment
This paper was supported in part by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2106961, Medium IIS-2403240, PIPP CCF-2200269), CDC MInD program, Meta faculty gift, and funds/computing resources from Georgia Tech and GTRI.
@@ Appendix A Related Works
Time-Series Forecasting. Recent works in deep learning have achieved notable achievements in time-series forecasting, such as RNNs, LSTNet, N-BEATS (Sherstinsky, 2020; Lai et al., 2018; Oreshkin et al., 2020). State-of-the-art models build upon the successes of self-attention mechanisms (Vaswani et al., 2017) with transformer-based architectures and significantly improve forecasting accuracy, such as Informer, Autoformer, Fedformer, PatchTST, iTransformer, FRNet (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Nie et al., 2023; Liu et al., 2024b; Zhang et al., 2024). However, these advanced models primarily rely on empirical risk minimization (ERM) with IID assumptions, i.e., train and test dataset follows the same data distribution, which exhibits limitations when potential distribution shifts in time series.
@@
General concept drift methods (via invariant learning) (Arjovsky et al., 2019; Ahuja et al., 2021; Krueger et al., 2021; Pezeshki et al., 2021; Sagawa et al., 2019) assume instances sampled from various environments and propose to identify and utilize invariant predictors across these environments. However, when applied to time-series forecasting, these methods encounter limitations. Additional methods specifically tailored for time series data also encounter certain constraints: DIVERSITY (Lu et al., 2023) is designed for time series classification and detection only. OneNet (Wen et al., 2024) is tailored solely for online forecasting scenarios using online ensembling. PeTS (Zhao et al., 2023) focuses on distribution shifts induced by the specific phenomenon of performativity.
Other works specifically tackle temporal shift issues in time-series forecasting (Kim et al., 2021; Liu et al., 2022; Fan et al., 2023; Liu et al., 2023). These approaches implement carefully crafted normalization strategies to ensure that both the lookback and horizon of a univariate time series adhere to normalized distributions. This alignment helps alleviate potential temporal shifts, where the statistical properties of the lookback and horizon time series may differ, over time.
@@ C.1 Datasets
We follow (Wu et al., 2022; Nie et al., 2023; Liu et al., 2024b) to preprocess data, which guides splitting datasets into train/validation/test sets and selecting the target variables. All datasets are preprocessed using the zero-mean normalization method.
@@ C.2 Baseline Implementation
We follow the commonly adopted setup for defining the forecasting horizon window length, as outlined in prior works (Wu et al., 2022; Nie et al., 2023; Liu et al., 2024b). Specifically, for datasets such as ETT and Exchange, the forecasting horizon windows are chosen from the set [96, 192, 336, 720], with a fixed lookback window size of 96 and a consistent label window size of 48 for the decoder (if required). Similarly, for the weekly reported ILI dataset, we employ forecasting horizon windows from [24, 36, 48, 60], with a fixed lookback window size of 36 and a constant label window size of 18 for the decoder (if required).

### v2

@@ 1 Introduction
Time-series forecasting finds applications in various real-world scenarios such as economics, urban computing, and epidemiology (Zhu and Shasha, 2002; Zheng et al., 2014; Deb et al., 2017; Mathis et al., 2024). These applications involve predicting future trends or events based on historical time-series data. For example, economists use forecasts to make financial and marketing plans, while sociologists use them to allocate resources and formulate policies for traffic or disease control.
@@ 2.2 Distribution Shift in Time Series
Temporal shift (also known as virtual shift (Tsymbal, 2004)) refers to phenomenon that the marginal probability distributions change over time, while the conditional distributions are the same.
@@
Concept drift (also known as real concept drift (Gama et al., 2014)111(Gama et al., 2014) defines concept drift as both virtual shift and real concept drift. Our concept drift definition is consistent with the definition of real concept drift in (Gama et al., 2014).) is the phenomenon where the conditional distributions change over time, while the marginal probability distributions are the same.
@@ 4.1 Setup
Datasets. We conduct experiments using six time-series datasets as leveraged in (Liu et al., 2024a): The daily reported currency exchange rates (Exchange) (Lai et al., 2018); The weekly reported influenza-like illness patients (ILI) (Kamarthi et al., 2021); Two-hourly/minutely reported electricity transformer temperature (ETTh1/ETTh2 and ETTm1/ETTm2, respectively) (Zhou et al., 2021). We follow the established experimental setups and target variable selections in previous works(Wu et al., 2021; 2022; Nie et al., 2023; Liu et al., 2024d). Datasets such as Traffic (PeMS) (Zhao et al., 2017) and Weather (Wu et al., 2021) are excluded from our evaluations, as their time series exhibit near-stationary behavior, with only moderate distribution shift issues. Further details on the dataset differences are discussed in Appendix C.1.
@@
Forecasting Model Baselines: ShifTS is model-agnostic, we include six time-series forecasting models (referred to as ‘Model’ in Table 1 and 4), including: Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), Crossformer (Zhang and Yan, 2022), PatchTST (Nie et al., 2023), TimeMixer (Wang et al., 2024) and iTransformer (Liu et al., 2024d), which of the last two are the state-of-the-art (SOTA) forecasting model. These models are used to demonstrate that ShifTS consistently enhances forecasting accuracy across various models, including SOTA.
@@ 4.2 Performance Improvement across Base Forecasting Models
| Dataset | ILI | Exchange | ETTh1 | ETTh2 | |||||
| Method | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| Base | ERM | 3.705 | 1.704 | 0.819 | 0.732 | 0.254 | 0.427 | 0.937 | 0.828 |
| Concept Drift Method | GroupDRO | 2.285 | 1.287 | 0.821 | 0.751 | 0.278 | 0.453 | 1.150 | 0.936 |
| IRM | 2.248 | 1.237 | 0.846 | 0.754 | 0.201 | 0.367 | 0.878 | 0.792 | |
| VREx | 2.285 | 1.286 | 0.821 | 0.742 | 0.314 | 0.486 | 1.142 | 0.938 | |
| EIIL | 2.036 | 1.159 | 0.822 | 0.749 | 0.212 | 0.433 | 1.122 | 0.930 | |
| Temporal Shift Method | RevIN | 0.815 | 0.708 | 0.475 | 0.476 | 0.085 | 0.224 | 0.205 | 0.358 |
| N-S Trans. | 0.781 | 0.688 | 0.484 | 0.481 | 0.086 | 0.226 | 0.203 | 0.355 | |
| SAN | 0.757 | 0.715 | 0.415 | 0.453 | 0.088 | 0.225 | 0.199 | 0.348 | |
| Combined Method | IRM+RevIN | 0.809 | 0.711 | 0.481 | 0.476 | 0.089 | 0.231 | 0.202 | 0.362 |
| EIIL+RevIN | 0.799 | 0.706 | 0.483 | 0.485 | 0.085 | 0.225 | 0.218 | 0.380 | |
| FOIL | 0.735 | 0.651 | 0.497 | 0.481 | 0.081 | 0.219 | 0.206 | 0.357 | |
| ShifTS (Ours) | 0.668 | 0.613 | 0.470 | 0.468 | 0.076 | 0.214 | 0.194 | 0.348 | |
@@ 4.3 Comparison with Distribution Shift Methods
The results highlight the advantages of ShifTS over existing distribution shift methods, achieving the highest average forecasting accuracy in 6 out of 8 evaluations, with the remaining 2 evaluations ranking second. Notably, as discussed in Section 3.2, we choose to use RevIN as it is one of the most popular yet simple and effective temporal shift methods.
@@
However, ShifTS is flexible and can integrate more advanced temporal shift methods to further enhance performance. While exploring these advanced temporal shift methods is beyond the scope of this work, we illustrate the potential benefits of such integration. For example, on the Exchange dataset, where SAN outperforms ShifTS, incorporating SAN in place of RevIN within ShifTS leads to even greater accuracy improvements. Detailed MSE values for these evaluations are provided in Table 3. Furthermore, the results underscore the importance of addressing concept drift using SAM when temporal shifts are effectively addressed.
@@ 6 Acknowledgment
This work was partly supported by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2403240, Medium IIS-2106961), NIH (1R01HL184139), CDC MInD program, Meta, and Dolby faculty gifts.
@@ Appendix A Related Works
Time-Series Forecasting. Recent works in deep learning have achieved notable achievements in time-series forecasting, such as RNNs, LSTNet, N-BEATS (Sherstinsky, 2020; Lai et al., 2018; Oreshkin et al., 2020). State-of-the-art models build upon the successes of self-attention mechanisms (Vaswani et al., 2017) with transformer-based architectures and significantly improve forecasting accuracy, such as Informer, Autoformer, Fedformer, PatchTST, iTransformer, FRNet (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Nie et al., 2023; Liu et al., 2024d; Zhang et al., 2024; Zhao et al., 2025c). However, these advanced models primarily rely on empirical risk minimization (ERM) with IID assumptions, i.e., train and test dataset follows the same data distribution, which exhibits limitations when potential distribution shifts in time series (Zhao et al., 2025b).
@@
General concept drift methods (via invariant learning) (Arjovsky et al., 2019; Ahuja et al., 2021; Krueger et al., 2021; Pezeshki et al., 2021; Sagawa et al., 2019) assume instances sampled from various environments and propose to identify and utilize invariant predictors across these environments. However, when applied to time-series forecasting, these methods encounter limitations. Additional methods specifically tailored for time series data also encounter certain constraints: DIVERSITY (Lu et al., 2023) is designed for time series classification and detection only. OneNet (Wen et al., 2024) is tailored solely for online forecasting scenarios using online ensembling. PeTS (Zhao et al., 2025a) focuses on distribution shifts induced by the specific phenomenon of performativity.
Other works explicitly address temporal distribution shift in time-series forecasting (Kim et al., 2021; Liu et al., 2022; Fan et al., 2023; Liu et al., 2023). These methods typically design normalization schemes that align the statistical properties of the lookback and forecast horizon, ensuring that both segments follow comparable normalized distributions. Such alignment mitigates temporal shift arising from discrepancies between historical observations and future targets. More recent time-series foundation models Aksu et al. (2024); Woo et al. (2024); Das et al. (2024); Liu et al. (2025b); Ansari et al. (2025) improve generalization by scaling model capacity and training on large and diverse corpora. Through large-scale pretraining and parameter scaling, these approaches may implicitly enhance robustness to distribution shift. However, most existing foundation models focus primarily on univariate settings and do not explicitly leverage rich exogenous information. In parallel, several recent studies (Liu et al., 2024b; c; ; 2025a; 2026) incorporate auxiliary textual context to assist time-series understanding and forecasting. These approaches extend beyond purely numerical modeling and explore multimodal conditioning, which is orthogonal to the scope of this work.
@@ C.1 Datasets
We follow (Wu et al., 2022; Nie et al., 2023; Liu et al., 2024d) to preprocess data, which guides splitting datasets into train/validation/test sets and selecting the target variables. All datasets are preprocessed using the zero-mean normalization method.
@@ C.2 Baseline Implementation
We follow the commonly adopted setup for defining the forecasting horizon window length, as outlined in prior works (Wu et al., 2022; Nie et al., 2023; Liu et al., 2024d). Specifically, for datasets such as ETT and Exchange, the forecasting horizon windows are chosen from the set [96, 192, 336, 720], with a fixed lookback window size of 96 and a consistent label window size of 48 for the decoder (if required). Similarly, for the weekly reported ILI dataset, we employ forecasting horizon windows from [24, 36, 48, 60], with a fixed lookback window size of 36 and a constant label window size of 18 for the decoder (if required).